Ten years from now, A will be twicw as old as B. Five years ago, A was three times as old as B. What are the present ages of A and B?
Answer by stanbon(75887)

Ten years from now, A will be twice as old as B. Five years ago, A was three times as old as B. What are the present ages of A and B?
----
Equations:
(A+10) = 2(B+10)
(A-5) = 3(B-5)
------
Rearrange:
A - 2B = 10
A - 3B = -10
----
Subtract and solve for "B":
B = 20
----
Solve for "A":
A = 2B+10
A = 2*20+10
A = 50
==================
cheers,
Stan H.